a boy is 24years younger than his father after two years their sum of ages will be 40. what is the age of fater?

Answer by ankor@dixie-net.com(22740)

boy is 24 yrs younger than his father
after two years their sum of ages will be 40.
what is the age of father?
:
Let f = age of father
Then
(f-24) = age of son
:
"after two years their sum of ages will be 40. "
(f+2) + (f-24) + 2 = 40
2f - 20 = 40
2f = 40 + 20
2f = 60
f = 30 yrs